# From multitype branching Brownian motions to branching Markov additive processes

### v1

@@ 2.3 Measure change by the additive martingale
Let $\mathbb{T}$ be the space of planar trees. We use the notation $v\prec u$ to denote that $v$ is an ancestor of $u$.
@@
$\widetilde{\mathbf{P}}^{\theta}_{x,i}[\xi_{t}=u;B]$ $=\frac{1}{e^{-\theta x}V_{i}(\theta)}\widetilde{\mathbf{E}}_{x,i}\left[\prod_{v\prec u}A_{v}e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbbm{1}_{B}\right]$ (2.25)
$=\frac{1}{e^{-\theta x}V_{i}(\theta)}\mathbf{E}_{x,i}\left[e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbbm{1}_{B}\right].$ (2.26)
@@
$\widetilde{\mathbf{E}}^{\theta}_{x,i}\left[\frac{e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbbm{1}_{B}}{W_{\theta}(t)}\right]=\frac{1}{e^{-\theta x}V_{i}(\theta)}\mathbf{E}_{x,i}\left[e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbbm{1}_{B}\right].$
@@
To describe the particle system under the law $\widetilde{\mathbf{P}}^{\theta}_{x,i}$, we introduce the following change of measure for MAPs, which is a variation of [PalRol2002, Proposition 5.6]; its proof is postponed to Section 5.
@@
$\widetilde{q}_{kj}=\frac{q_{kj}V_{j}(\theta)G_{kj}(\theta)}{V_{k}(\theta)},\;\quad\mathbb{P}(\widetilde{U}_{kj}\in\mathrm{d}x)=\frac{e^{-\theta x}}{G_{kj}(\theta)}\mathbb{P}(U_{kj}\in\mathrm{d}x),\quad\forall j\neq k,$ (2.29)
$\widetilde{q}_{kk}=-\sum_{j\neq k}\widetilde{q}_{kj},\;\widetilde{\sigma}_{k}=\sigma_{k},\;\widetilde{\mathtt{a}}_{k}=\mathtt{a}_{k}-\theta\sigma_{k}^{2}-\int x\mathbbm{1}_{\left\{[0,1]\right\}}(|x|)(1-e^{-\theta x})\Lambda_{k}(\mathrm{d}x),\;\widetilde{\Lambda}_{k}(\mathrm{d}x)=e^{-\theta x}\Lambda_{k}(\mathrm{d}x).$
Therefore the corresponding MAP triplet is given by $\left((\widetilde{\phi}_{i})_{i\in\mathcal{I}},\widetilde{Q}:=(\widetilde{q}_{kj})_{k,j\in\mathcal{I}},\widetilde{G}:=(\widetilde{G}_{kj})_{k,j\in\mathcal{I}}\right)$, where $\widetilde{\phi}_{k}(\alpha)=\phi_{k}(\alpha+\theta)-\phi_{k}(\theta)$ for $k\in\mathcal{I}$, and $\widetilde{G}_{kj}(\alpha)=\frac{G_{kj}(\alpha+\theta)}{G_{kj}(\theta)}$ for $k,j\in\mathcal{I}$.
@@
  • Given the type process $J_{\xi}(t)$ of the spine, the branching rate of the spine at time $t\geq 0$ is given by $\beta_{J_{\xi}(t)}m_{J_{\xi}(t)}$; when it splits, it gives birth to an offspring of the same type at the same position. The number of children is given by the size-biased law $(\widetilde{\mu}_{j}(k):=\frac{k\mu_{j}(k)}{m_{j}},k\geq 1)$, for $j\in\mathcal{I}$.
@@
$\frac{e^{\theta x}}{V_{i}(\theta)}\mathbf{E}_{x,i}\bigg[\sum_{u\in\mathcal{N}_{t}}g(X_{u}(t),J_{u}(t))e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\bigg]=\widetilde{\mathbf{E}}^{\theta}_{x,i}[g(X_{\xi}(t),J_{\xi}(t))]=\mathbb{E}^{\theta}_{x,i}[g(\chi_{t},\Theta_{t})].$ (2.30)
@@
Before proving Theorem 1.4, we first give some lemmas.
@@
Suppose that at least one of the following two conditions hold: (i) $\theta\lambda^{\prime}(\theta)\geq\lambda(\theta)$; (ii) there exists some $j\in\mathcal{I}$ such that $\sum_{k\geq 1}(k\log k)\mu_{j}(k)=+\infty$. We next show that $\limsup_{t\to\infty}W_{\theta}(t)=+\infty$ $\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s. and then by [PTEDurrett, Theorem 4.3.5] we conclude that $W_{\theta}(t)$ converges to 0, $\mathbf{P}_{x,i}$-a.s.
@@
When $\theta\lambda^{\prime}(\theta)=\lambda(\theta)$, by Proposition 2.1, we also have (2.32). Therefore $\limsup_{t\to\infty}W_{\theta}(t)=+\infty,\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s.
@@
Let $N_{j}(n)$ denote the total number of fissions (branchings) in the spine $\xi$ (across all types) by the time state $j$ undergoes its $n$-th fission. Furthermore, let $T_{n}$ be the $n$-th fission time of the spine. Under this definition, at time $T_{N_{j}(n)}$, the spine $\xi$ undergoes its $n$-th fission on state $j$ for all $n\geq 1$. Therefore $\{A_{\xi}(T_{N_{j}(n)}):=A_{\xi_{T_{N_{j}(n)}}}\}_{n\geq 1}$ is an i.i.d. sequence with law $(\widetilde{\mu}_{j}(k))_{k\geq 1}$. Since ${\widetilde{\mathbf{E}}_{x,i}}^{\theta}\left[\log A_{\xi}(T_{N_{j}(n)})\right]=\frac{1}{m_{j}}\sum_{k\geq 1}(k\log k)\mu_{j}(k)=+\infty$, by Borel-Cantelli’s lemma,
@@
$\limsup_{n\to\infty}\frac{\log A_{\xi}(T_{N_{j}(n)})}{n}=+\infty,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}.$ (2.33)
For $t\geq 0$, $j\in\mathcal{I}$, define $D_{j}(t):=\int_{0}^{t}\mathbbm{1}_{\left\{J_{\xi}(s)=j\right\}}\mathrm{d}s$ as the cumulative time the spine spends in state $j$ by time $t$. For $n\geq 1$, $D_{j}(T_{N_{j}(n)})$ denotes the cumulative time spent in state $j$ up to the occurrence of the $n$-th fission event on state $j$. Since the spine undergoes fissions at rate $\widetilde{\beta}_{j}:=\beta_{j}m_{j}$ when in state $j$, the strong law of large numbers implies
$\lim_{n\to\infty}\frac{D_{j}(T_{N_{j}(n)})}{n}=\frac{1}{\widetilde{\beta}_{j}},\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.34)
Since $(J_{\xi}(t),t\geq 0)$ is an irreducible Markov chain under $\widetilde{\mathbf{P}}^{\theta}_{x,i}$, we have $\lim_{t\to\infty}\frac{D_{j}(t)}{t}=\widetilde{\pi}_{j}$ a.s., where $(\widetilde{\pi}_{j})_{j\in\mathcal{I}}$ is the invariant distribution of $(J_{\xi}(t),t\geq 0)$. Notice that $T_{N_{j}(n)}\rightarrow\infty$ as $n\rightarrow\infty$, then we have
$\lim_{n\to\infty}\frac{D_{j}(T_{N_{j}(n)})}{T_{N_{j}(n)}}=\widetilde{\pi}_{j},\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.35)
@@
According to Theorem 2.7, the process $(T_{n})_{n\geq 1}$ is a Cox process on $\mathbb{R}_{+}$ directed by $({J}_{\xi}(t))_{t\geq 0}$ with rate $\widetilde{\beta}_{{J}_{\xi}(t)}\mathrm{d}t$. Since $\min_{l\in\mathcal{I}}\widetilde{\beta}_{l}\leq\widetilde{\beta}_{{J}_{\xi}(t)}\leq\max_{l\in\mathcal{I}}\widetilde{\beta}_{l}$ for all $t\geq 0$, the $n$-th arrival time $T_{n}$ is stochastically bounded by the $n$-th arrival time of two Poisson point processes with intensity $\max_{l\in\mathcal{I}}\widetilde{\beta}_{l}\mathrm{d}t$ and $\min_{l\in\mathcal{I}}\widetilde{\beta}_{l}\mathrm{d}t$. Then we have
@@
$\liminf_{n\to\infty}\frac{T_{N_{j}(n)}}{N_{j}(n)}\geq\liminf_{n\to\infty}\frac{T_{n}}{n}\geq\frac{1}{\max_{l\in\mathcal{I}}\widetilde{\beta}_{l}},\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.36)
@@
$\limsup_{n\to\infty}\frac{T_{n}}{n}\leq\frac{1}{\min_{l\in\mathcal{I}}\widetilde{\beta}_{l}},\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.37)
Combining (2.34), (2.35), and (2.36) we have
@@
$\liminf_{n\to\infty}\frac{n}{N_{j}(n)}=\liminf_{n\to\infty}\frac{n}{D_{j}(T_{N_{j}(n)})}\frac{D_{j}(T_{N_{j}(n)})}{T_{N_{j}(n)}}\frac{T_{N_{j}(n)}}{N_{j}(n)}\geq\frac{\widetilde{\beta}_{j}\widetilde{\pi}_{j}}{\max_{l\in\mathcal{I}}\widetilde{\beta}_{l}}>0,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.38)
By (2.38) and (2.33), we have
$\limsup_{n\to\infty}\frac{\log A_{\xi}(T_{N_{j}(n)})}{N_{j}(n)}\geq\limsup_{n\to\infty}\frac{\log A_{\xi}(T_{N_{j}(n)})}{n}\liminf_{n\to\infty}\frac{n}{N_{j}(n)}=+\infty\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.},$
which implies
@@
$\limsup_{n\to\infty}\frac{\log A_{\xi}(T_{n})}{n}=+\infty\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.39)
At time $T_{n}$, we have a lower bound
$W_{\theta}(T_{n})$ $\geq A_{\xi}(T_{n})e^{-\theta X_{\xi}(T_{n})-\lambda(\theta)T_{n}}V_{J_{\xi}(T_{n})}(\theta)$ (2.40)
$=\exp\left\{n\left(\frac{\log A_{\xi}(T_{n})}{n}-\frac{T_{n}}{n}\left(\frac{\theta X_{\xi}(T_{n})}{T_{n}}+\lambda(\theta)\right)\right)\right\}V_{J_{\xi}(T_{n})}(\theta).$ (2.41)
@@
In the case of $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\sum_{k\geq 1}(k\log k)\mu_{j}(k)=+\infty$ for some $j\in\mathcal{I}$, by (2.37), (2.39) and (2.41), we again have $\limsup_{t\to\infty}W_{\theta}(t)=+\infty,\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s.
@@
On the other hand, assume that $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\sum_{k\geq 1}(k\log k)\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$. We prove that the additive martingale converges in $\mathcal{L}^{1}(\mathbf{P}_{x,i})$. Recall that $\mathcal{G}$ is defined by (2.19), being the $\sigma$-field generated by the motion and type of the spine, its fission time and its children. By [ShiBRW, Lemma 4.2], it suffices to show:
@@
By (2.43), we have
$\widetilde{\mathbf{E}}_{x,i}^{\theta}\left[W_{\theta}(t)\mid\mathcal{G}\right]$ $=e^{-\theta X_{\xi}(t)-\lambda(\theta)t}V_{J_{\xi}(t)}(\theta)+\sum_{n=1}^{\infty}\mathbbm{1}_{\left\{T_{n}\leq t\right\}}A_{\xi}(T_{n})e^{-\theta X_{\xi}(T_{n})-\lambda(\theta)T_{n}}V_{J_{\xi}(T_{n})}(\theta)$
$=e^{-\theta X_{\xi}(t)-\lambda(\theta)t}V_{J_{\xi}(t)}(\theta)+\sum_{j\in\mathcal{I}}\sum_{n=1}^{\infty}\mathbbm{1}_{\left\{T_{n}\leq t\right\}}A_{\xi}(T_{N_{j}(n)})e^{-\theta X_{\xi}({T_{N_{j}(n)}})-\lambda(\theta)T_{N_{j}(n)}}V_{j}(\theta)$
$=e^{-\theta X_{\xi}(t)-\lambda(\theta)t}V_{J_{\xi}(t)}(\theta)$
$\quad+\sum_{j\in\mathcal{I}}\sum_{n=1}^{\infty}\mathbbm{1}_{\left\{T_{n}\leq t\right\}}\exp{\left\{N_{j}(n)\left[\frac{\log A_{\xi}(T_{N_{j}(n)})}{N_{j}(n)}+\frac{T_{N_{j}(n)}}{N_{j}(n)}\left(-\frac{\theta X_{\xi}({T_{N_{j}(n)}})}{T_{N_{j}(n)}}-\lambda(\theta)\right)\right]\right\}}V_{j}(\theta).$
When $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$, the first term above converges to 0, $\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s. Since $\sum_{k\geq 1}(k\log k)\mu_{j}(k)<+\infty$ for $\forall j\in\mathcal{I}$, by the Borel-Cantelli lemma, we have
@@
Therefore, for any $j\in\mathcal{I}$, by (2.36), there exists $c_{j}>0$ such that
$\limsup_{n\to\infty}\left(\frac{\log A_{\xi}(T_{N_{j}(n)})}{N_{j}(n)}+\frac{T_{N_{j}(n)}}{N_{j}(n)}\left(-\frac{\theta X_{\xi}(T_{N_{j}(n)})}{T_{N_{j}(n)}}-\lambda(\theta)\right)\right)\leq c_{j}\left(\theta\lambda^{\prime}(\theta)-\lambda(\theta)\right)<0,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\mbox{-a.s.}$
@@
Define $w(i):=\mathbf{P}_{x,i}(W_{\theta}(\infty)=0)$, for $i\in\mathcal{I}$, and $\vec{w}:=(w(1),\ldots,w(\mathtt{d}))$. Note that $w(i)$ does not depend on $x$. By [PTEDurrett, Theorem 4.3.5], when $W_{\theta}$ converges in $\mathcal{L}^{1}(\mathbf{P}_{x,i})$, we have $\mathbf{E}_{x,i}\left[\frac{W_{\theta}(\infty)}{W_{\theta}(0)}\right]=1$. This implies $\vec{w}\neq(1,\dots,1)$. By strong Markov property on the first branching/type-changing time, we have

### v2

@@ 2.3 Measure change by the additive martingale
Let $\mathbb{T}$ be the space of planar trees. We denote $v\preceq u$ if $v$ is an ancestor of $u$. Write $v\prec u$ if $v\preceq u$ and $v\neq u$.
@@
$\widetilde{\mathbf{P}}^{\theta}_{x,i}[\xi_{t}=u;B]$ $=\frac{1}{e^{-\theta x}V_{i}(\theta)}\widetilde{\mathbf{E}}_{x,i}\left[\prod_{v\prec u}A_{v}e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbb{1}_{B}\right]$ (2.30)
$=\frac{1}{e^{-\theta x}V_{i}(\theta)}\mathbf{E}_{x,i}\left[e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbb{1}_{B}\right].$ (2.31)
@@
$\widetilde{\mathbf{E}}^{\theta}_{x,i}\left[\frac{e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbb{1}_{B}}{W_{\theta}(t)}\right]=\frac{\mathbf{E}_{x,i}\left[e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{u\in\mathcal{N}_{t}\right\}}\mathbb{1}_{B}\right]}{e^{-\theta x}V_{i}(\theta)}.$ (2.32)
@@
To describe the particle system under the law $\widetilde{\mathbf{P}}^{\theta}_{x,i}$, we introduce the following change of measure for MAPs, which is a variation of [47, Proposition 5.6]; its proof is postponed to Section 5.
@@
$\widetilde{q}_{kj}=\frac{q_{kj}V_{j}(\theta)G_{kj}(\theta)}{V_{k}(\theta)},\;\quad\mathbb{P}(\widetilde{U}_{kj}\in\mathrm{d}x)=\frac{e^{-\theta x}}{G_{kj}(\theta)}\mathbb{P}(U_{kj}\in\mathrm{d}x),\quad\forall j\neq k,\quad\widetilde{q}_{kk}=-\sum_{j\neq k}\widetilde{q}_{kj},$ (2.35)
$\widetilde{\sigma}_{k}=\sigma_{k},\;\widetilde{\mathtt{a}}_{k}=\mathtt{a}_{k}-\theta\sigma_{k}^{2}-\int x\mathbbm{1}_{\left\{[0,1]\right\}}(|x|)(1-e^{-\theta x})\Lambda_{k}(\mathrm{d}x),\;\widetilde{\Lambda}_{k}(\mathrm{d}x)=e^{-\theta x}\Lambda_{k}(\mathrm{d}x).$
Therefore the corresponding MAP triplet is given by
$\left((\widetilde{\phi}_{k})_{k\in\mathcal{I}},\widetilde{Q}:=(\widetilde{q}_{kj})_{k,j\in\mathcal{I}},\widetilde{G}:=(\widetilde{G}_{kj})_{k,j\in\mathcal{I}}\right),$ (2.36)
where $\widetilde{\phi}_{k}(\alpha)=\phi_{k}(\alpha+\theta)-\phi_{k}(\theta)$ for $k\in\mathcal{I}$, and $\widetilde{G}_{kj}(\alpha)=\frac{G_{kj}(\alpha+\theta)}{G_{kj}(\theta)}$ for $k,j\in\mathcal{I}$.
@@
  • Given the type process $J_{\xi}(t)$ of the spine, the branching rate of the spine at time $t\geq 0$ is given by $\widetilde{\beta}_{J_{\xi}(t)}$, where $\widetilde{\beta}_{i}=\beta_{i}m_{i},i\in\mathcal{I}$; when it splits, it gives birth to an offspring of the same type at the same position. The number of children is given by the size-biased law $(\widetilde{\mu}_{j}(k):=\frac{k\mu_{j}(k)}{m_{j}},k\geq 1)$, for $j\in\mathcal{I}$ with $m_{j}>0$; if $m_{j}=0$, then $\widetilde{\beta}_{j}=0$, meaning that the spine can never split at state $j$, and we simply define by convention that $\widetilde{\mu}_{j}(1)=1$.
@@
$\frac{e^{\theta x}}{V_{i}(\theta)}\mathbf{E}_{x,i}\bigg[\sum_{u\in\mathcal{N}_{t}}g(X_{u}(t),J_{u}(t))e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\bigg]$ (2.37)
$=$ $\,\widetilde{\mathbf{E}}^{\theta}_{x,i}[g(X_{\xi}(t),J_{\xi}(t))]=\mathbb{E}^{\theta}_{x,i}[g(\chi_{t},\Theta_{t})].$ (2.38)
@@
Suppose that at least one of the following two conditions hold: (i) $\theta\lambda^{\prime}(\theta)\geq\lambda(\theta)$; (ii) there exists some $j\in\mathcal{I}$ such that $\sum_{k\geq 1}(k\log k)\mu_{j}(k)=+\infty$ (then we must have $\beta_{j}m_{j}>0$). We next show that in either case $\limsup_{t\to\infty}W_{\theta}(t)=+\infty$ $\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s. and then by [20, Theorem 4.3.5] we conclude that $W_{\theta}(t)$ converges to 0, $\mathbf{P}_{x,i}$-a.s.
@@
When $\theta\lambda^{\prime}(\theta)=\lambda(\theta)$, by Proposition 2.1, we also have (2.40). Therefore, we have
$\limsup_{t\to\infty}W_{\theta}(t)=+\infty,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\mbox{-a.s.}$ (2.41)
@@
Let $T_{n}$ denote the $n$-th fission (branching) time of the spine. For state $j$, let $T_{m}^{j}$ be the time the spine undergoes its $m$-th fission at state $j$. Therefore $\{A_{\xi}(T_{n}^{j}):=A_{\xi_{T_{n}^{j}}}\}_{n\geq 1}$ is an i.i.d. sequence with law $(\widetilde{\mu}_{j}(k))_{k\geq 1}$. Since ${\widetilde{\mathbf{E}}_{x,i}}^{\theta}\left[\log A_{\xi}(T_{n}^{j})\right]=\frac{1}{m_{j}}\sum_{k\geq 1}(k\log k)\mu_{j}(k)=\infty$, the Borel-Cantelli lemma leads to
@@
$\limsup_{n\to\infty}\frac{\log A_{\xi}(T_{n}^{j})}{n}=+\infty,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.42)
@@
According to Theorem 2.7, the process $(T_{m}^{j})_{m\geq 1}$ is a Cox process on $\mathbb{R}_{+}$ with rate $\widetilde{\beta}_{j}\mathbbm{1}_{\left\{{J}_{\xi}(t)=j\right\}}\mathrm{d}t$, directed by $({J}_{\xi}(t))_{t\geq 0}$, where $\widetilde{\beta}_{j}:=\beta_{j}m_{j}$ and $(J_{\xi}(t),t\geq 0)$ is an irreducible Markov chain specified by (2.35). Let $(\widetilde{\pi}_{j})_{j\in\mathcal{I}}$ be the invariant distribution of $(J_{\xi}(t),t\geq 0)$. Then we have
@@
$\lim_{n\to\infty}\frac{T_{n}^{j}}{n}$ $=\frac{1}{\widetilde{\pi}_{j}\widetilde{\beta}_{j}},\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.43)
Let us justify (2.43). We have by the ergodic theorem that
@@
$\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\widetilde{\beta}_{j}\mathbbm{1}_{\left\{{J}_{\xi}(s)=j\right\}}\mathrm{d}s=\widetilde{\pi}_{j}\widetilde{\beta}_{j},\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.44)
Given $J_{\xi}$, the counting process $(N^{j}(t):=\max\{n:T_{n}^{j}\leq t\})_{t\geq 0}$ is an inhomogeneous Poisson process with rate $\widetilde{\beta}_{j}\mathbbm{1}_{\left\{{J}_{\xi}(t)=j\right\}}\mathrm{d}t$. Notice that as $t\to\infty$, the integral
@@
$\int_{0}^{t}\widetilde{\beta}_{J_{\xi}(s)}\mathbbm{1}_{\left\{{J}_{\xi}(s)=j\right\}}\mathrm{d}s\to\infty,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\mbox{-a.s.}$ (2.45)
Therefore, we have
$\widetilde{\mathbf{P}}_{x,i}^{\theta}\left(\lim_{t\to\infty}\frac{N^{j}(t)}{\int_{0}^{t}\widetilde{\beta}_{j}\mathbbm{1}_{\left\{{J}_{\xi}(s)=j\right\}}\mathrm{d}s}=1\Bigg|J_{\xi}\right)=1.$
It follows that
@@
$\lim_{t\to\infty}\frac{N^{j}(t)}{\int_{0}^{t}\widetilde{\beta}_{j}\mathbbm{1}_{\left\{{J}_{\xi}(s)=j\right\}}\mathrm{d}s}=1,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$ (2.46)
Combining (2.44) and (2.46) we have
$\lim_{t\to\infty}\frac{N^{j}(t)}{t}=\widetilde{\pi}_{j}\widetilde{\beta}_{j}>0,\quad\widetilde{\mathbf{P}}_{x,i}^{\theta}\text{-a.s.}$
Taking $t=T_{n}^{j}$ and noticing that $N^{j}(T_{n}^{j})=n$, we deduce (2.43).
At time $T_{n}^{j}$, we have a lower bound of the additive martingale,
$W_{\theta}(T_{n}^{j})$ $\geq A_{\xi}(T_{n}^{j})e^{-\theta X_{\xi}(T_{n}^{j})-\lambda(\theta)T_{n}^{j}}V_{j}(\theta)$ (2.47)
$=\exp\left\{n\left(\frac{\log A_{\xi}(T_{n}^{j})}{n}-\frac{T^{j}_{n}}{n}\left(\frac{\theta X_{\xi}(T_{n}^{j})}{T_{n}^{j}}+\lambda(\theta)\right)\right)\right\}V_{j}(\theta).$ (2.48)
@@
In the case of $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\sum_{k\geq 1}(k\log k)\mu_{j}(k)=+\infty$ for $j\in\mathcal{I}$ with $\beta_{j}m_{j}>0$, by (2.42), (2.43) and (2.48), we again have $\limsup_{t\to\infty}W_{\theta}(t)=+\infty,\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s.
@@
On the other hand, assume that $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$ and $\sum_{k\geq 1}(k\log k)\mu_{j}(k)<\infty$ for all $j\in\mathcal{I}$. We prove that the additive martingale converges in $\mathcal{L}^{1}(\mathbf{P}_{x,i})$. Recall that $\mathcal{G}$ defined by (2.24) is the $\sigma$-field generated by the motion and type of the spine with its fission time and children. By [54, Lemma 4.2], it suffices to show:
@@
When $\theta\lambda^{\prime}(\theta)<\lambda(\theta)$, the first term above converges to 0, $\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s. For the second term, we have
$\sum_{n=1}^{\infty}\mathbbm{1}_{\left\{T_{n}\leq t\right\}}A_{\xi}(T_{n})e^{-\theta X_{\xi}(T_{n})-\lambda(\theta)T_{n}}V_{J_{\xi}(T_{n})}(\theta)$
$=\sum_{j\in\mathcal{I}}\sum_{n=1}^{\infty}\mathbbm{1}_{\left\{T_{n}\leq t\right\}}A_{\xi}(T_{N_{j}(n)})e^{-\theta X_{\xi}({T_{N_{j}(n)}})-\lambda(\theta)T_{N_{j}(n)}}V_{j}(\theta)$
$=\sum_{j\in\mathcal{I}}\sum_{n=1}^{\infty}\mathbbm{1}_{\left\{T_{n}\leq t\right\}}V_{j}(\theta)\exp{\left\{N_{j}(n)\left[\frac{\log A_{\xi}(T_{N_{j}(n)})}{N_{j}(n)}+\frac{T_{N_{j}(n)}}{N_{j}(n)}\left(-\frac{\theta X_{\xi}({T_{N_{j}(n)}})}{T_{N_{j}(n)}}-\lambda(\theta)\right)\right]\right\}}.$
Since $\sum_{k\geq 1}(k\log k)\mu_{j}(k)<+\infty$ for $\forall j\in\mathcal{I}$, by the Borel-Cantelli lemma, we have
@@
Therefore, for any $j\in\mathcal{I}$, by (2.43), there exists $c_{j}>0$ such that $\widetilde{\mathbf{P}}_{x,i}^{\theta}$-a.s.
$\limsup_{n\to\infty}\left(\frac{\log A_{\xi}(T_{N_{j}(n)})}{N_{j}(n)}+\frac{T_{N_{j}(n)}}{N_{j}(n)}\left(-\frac{\theta X_{\xi}(T_{N_{j}(n)})}{T_{N_{j}(n)}}-\lambda(\theta)\right)\right)\leq c_{j}\left(\theta\lambda^{\prime}(\theta)-\lambda(\theta)\right)<0.$
@@
Define $w(i):=\mathbf{P}_{x,i}(W_{\theta}(\infty)=0)$, for $i\in\mathcal{I}$, and $\vec{w}:=(w(1),\ldots,w(\mathtt{d}))$. Note that $w(i)$ does not depend on $x$. By [20, Theorem 4.3.5], when $W_{\theta}$ converges in $\mathcal{L}^{1}(\mathbf{P}_{x,i})$, we have $\mathbf{E}_{x,i}\left[\frac{W_{\theta}(\infty)}{W_{\theta}(0)}\right]=1$. This implies $\vec{w}\neq(1,\dots,1)$. By strong Markov property on the first branching/type-changing time, we have